OpenStudy (anonymous):

Solve each quadratic, leave answers as simplified radicals as necessary. 3x^2-13x=10

OpenStudy (anonymous):

Quadratic is ax^2+bx+c=0 So what do you do first?

OpenStudy (anonymous):

Well, set it equal to zero. So \[3x ^{2}-13x-10=0\]

OpenStudy (anonymous):

Yep..now solve using the quadratic formula.

OpenStudy (anonymous):

Alright, here is what I got: x=-2/3 and x=5

OpenStudy (anonymous):

Did you plug them into the equation to verify they work?

OpenStudy (anonymous):

I am right now, one moment.

OpenStudy (anonymous):

They both equal zero.

OpenStudy (anonymous):

But something tells me that I must try it with ten out of the modified equation.

OpenStudy (anonymous):

If they both equal 0 then they are right.

OpenStudy (anonymous):

Alright, that is one down, may I show you the next one?

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

Alright, for these I have a feeling my current answers are incorrect. \[2x ^{2}-8x=6\]

OpenStudy (anonymous):

Same thing we did above

OpenStudy (anonymous):

Here are my answers: \[x=-\sqrt{7}-2 \]

OpenStudy (anonymous):

I got a positive version of the above as well.

OpenStudy (anonymous):

How would I check my answer?

OpenStudy (anonymous):

Let's do this the easy way. First move the 6 and what does the equation look like?

OpenStudy (anonymous):

\[2x ^{2}-8x-6=0\]

OpenStudy (anonymous):

Now simplify that equation to get what?

OpenStudy (anonymous):

\[2(x^{2}-4x-3)\]

OpenStudy (anonymous):

I do have a feeling that I should've used the quadratic, but my gut led me with this.

OpenStudy (anonymous):

Actually just x^2-4x-3=0

OpenStudy (anonymous):

Where does the 2 go?

OpenStudy (anonymous):

We are simplifying not factoring. We just divided everything by 2.

OpenStudy (anonymous):

Alright. So, do I apply the quadratic still?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

But stop because you have the right answer above.

OpenStudy (anonymous):

\[2+/- \sqrt{7}\]

OpenStudy (anonymous):

Alright.

OpenStudy (anonymous):

Now this one I couldn't figure out how to do to save my life.

OpenStudy (anonymous):

2x^2+3x+4=0

OpenStudy (anonymous):

Plug into the formula

OpenStudy (anonymous):

\[-3\pm \sqrt{3^{2}-4(2)(3)}\div4\]

OpenStudy (anonymous):

That it?

OpenStudy (anonymous):

4(2)(4) I believe so what does that give you?

OpenStudy (anonymous):

So, 9-32.